OpenStudy (falling_in_katt):

***WILL MEDAL*** Divide (x ^ 2 - 4)(x - 1)

OpenStudy (anonymous):

"Simplified" using the Mathematica program. \[x-1+\frac{3}{x-1}\]

OpenStudy (anonymous):

Assuming that Divide (x ^ 2 - 4)(x - 1) -> (x ^ 2 - 4)/(x - 1)

OpenStudy (campbell_st):

I think there may be a typo in the answer above just look at the problem this way \[\frac{x^2 - 4}{x -1} = \frac{(x^2 - 1) - 3}{(x -1)}\] the fraction can be split \[\frac{x^2 -1}{x -1} - \frac{3}{x -1}\] a little factorising \[\frac{(x -1)(x + 1)}{(x -1)} - \frac{3}{(x -1)}\] cancel the common factor in the 1st fraction for the solution